OpenStudy (anonymous):

The area of a triangle is 24 square inches. What is the height of the triangle if the base length is 4 inches? 6 inches 8 inches 12 inches 20 inches

OpenStudy (anonymous):

@Vocaloid

Vocaloid (vocaloid):

hint: area = (1/2)*(height)*(base)

OpenStudy (anonymous):

so 1/2 * 24 * 4 ?

OpenStudy (anonymous):

@Vocaloid

Vocaloid (vocaloid):

read the problem carefully...

Vocaloid (vocaloid):

area = 24 base = 4 24 = (1/2)*(4)*(height) solve for height

Vocaloid (vocaloid):

start by simplifying.... (1/2)*4*height = 2*height 24 = 2*height height = ?

OpenStudy (anonymous):

4 sorry i had to go out @Vocaloid

Vocaloid (vocaloid):

not quite.... 24 = 2 times height divide both sides by 2

OpenStudy (anonymous):

12 and 1 @Vocaloid

OpenStudy (anonymous):

so is my answer 12?

Vocaloid (vocaloid):

right

OpenStudy (anonymous):

thank you! :)